OpenStudy (anonymous):

More math help D:

OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x – 4. What is f(x) • g(x)?

OpenStudy (anonymous):

Answers: x2 – 12x + 4 x2 + 4x + 4 x2 + 4x – 32 x2 – 4x – 32

OpenStudy (anonymous):

f(x) . g(x) = ( x + 8 )( x - 4 )= x^2 + 4x - 32 Simple algebra=c.

OpenStudy (jiteshmeghwal9):

f(x).g(x) put the values of f(x) & g(x) (x+8)(x-4)=x(x+8)+8(x-4)=?

OpenStudy (jiteshmeghwal9):

@anas2000 don't give direct answers since it is against the code of conduct :)

OpenStudy (anonymous):

Thank you

OpenStudy (jiteshmeghwal9):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

i didn't i also explained the question